monkica:

Eight less than four times a number is less than 56. What are the possible values of that number?

DuarteME:

Let \(n\) be the unknown number, which I'm assuming to be natural \((n \in \mathbb{N})\). We therefore know that: \(8 < 4n < 56.\) This double inequality is actually the conjunction of two simple inequalities: \(8 < 4n \quad \textrm{and}\quad 4n < 56.\) Can you proceed from now?

DuarteME:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @nicolegunter Any number less than 16 will work. \(\color{#0cbb34}{\text{End of Quote}}\) I don't think so. For instance \(4 \times 1 = 4\), which is not greater than \(8\) and \(15 \times 4 = 60\), which is not less than \(56\).

nicolegunter:

@DuarteME

DuarteME:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @RichardWatterson7 It would start off like 4x-8<56 (then you add 8 on both sides then it'll end up like 4x<64 (then you divide 4 on both sides) then you'll get x<16 \(\color{#0cbb34}{\text{End of Quote}}\) I see what you mean! However, I think that the question is regarding \(8 < 4x < 56\) rather than \(4x - 8 <56\).

nicolegunter:

ANY NUMBER LESS THE 16 WILL WORK

DuarteME:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @RichardWatterson7 2<x<14 \(\color{#0cbb34}{\text{End of Quote}}\) I agree with you! :)

RichardWatterson7:

👌 Gn 2 all nd talk 2 y'all 2morrow